# ISRE-$0$-$0$.00

### v2

@@ 4 Model
Motivated by our theoretical inquiry and exploratory analyses, we develop a Bayesian learning model to investigate the dynamics of individual adoption behaviors. Whenever a user would like to access resources from the Internet, she must decide whether to start the process with ChatGPT or traditional search engines. This decision is based on the perceived utility derived from each option, which evolves through a learning process influenced by the user’s prior experiences. Consequently, the final choice is determined through utility maximization. The utility function, learning process, and decision-making procedure will be elaborated upon in the following subsections.
@@ 4.1 Utility Function
Given that traditional search engines have been active for a long period and users are already familiar with this option, we assume that its utility $u_{it0}$ can be decomposed into a constant term $c$ and a stochastic error term $\varepsilon_{it0}$, in line with the Bayesian learning literature (Zheng et al. 2020):
$u_{it0}=c+\varepsilon_{it0}$ (1)
@@
As introduced in RQ2, users with different demographic backgrounds may derive varying levels of utility from using ChatGPT. Therefore, we model each individual’s demographic attributes as contributing linearly to her representative utility, allowing us to characterize the potential utility divide.
$v_{i}=$ $\alpha_{0i}+\alpha_{1}\cdot HighEdu_{i}+\alpha_{2}\cdot Age_{i}+\alpha_{3}\cdot Male_{i}+\alpha_{4}\cdot White_{i}+\alpha_{5}\cdot IT_{i}$ (3)
Note that in other discrete choice settings (Train 2009), only differences in utility are relevant; therefore, only one constant term is needed in the utility functions of the two alternatives to capture the baseline relative differences between ChatGPT and the traditional search engine. However, in our learning model, where we normalize the distribution of individuals’ initial beliefs, both constants can be identified, as discussed in Section 5. Additionally, the parameters $(\alpha_{1},\ldots,\alpha_{5})$ also capture differences in relative utility instead of absolute utility. For instance, a positive $\alpha_{1}$ would indicate that users with higher education tend to derive greater relative gains from using ChatGPT compared to traditional search engines.
However, unlike traditional search engines, ChatGPT has been launched only recently, and individuals may not initially be fully certain about ChatGPT’s utility $u_{it1}$. Therefore, before a user $i$ decides whether to use ChatGPT to fulfill her query, she first forms an expectation of the utility based on the information available up to period $t$ (hereinafter referred to as “perceived utility”), and then choose the alternative that maximizes her expected utility. The perception formation process follows the Bayesian learning framework (Wang and Yang 2024), and will be elucidated in Section 4.2.
@@ 4.2 Individuals’ Learning Process
Since traditional search engines have been available for an extended period, individuals are already familiar with them, and their beliefs are assumed to have stabilized prior to the observation period. Consequently, similar to other Bayesian learning models (Zheng et al. 2020), the perceived representative utility of the outside option $E[c|I_{it}]$ is assumed to be constant for consumers a priori and is not influenced by new information.
$E[u_{it0}|I_{it}]=E[c|I_{it}]+\varepsilon_{it0}=c+\varepsilon_{it0}$ (6)
In contrast, the utility of the newly introduced ChatGPT is less certain to users, resulting in a dynamic update of their belief $E[u_{it1}|I_{it}]$ over time as they gather more information and experience with the tool. Following the Bayesian learning literature (Wang and Yang 2024, Erdem and Keane 1996), we assume that prior to any direct experience with ChatGPT, individuals possess some prior information regarding the representative utility of using ChatGPT, which is modeled as following a normal distribution $N(v_{0},\sigma_{0}^{2})$. The individual perceives that the actual representative utility $v_{i}$ is drawn from this distribution, where the mean of this distribution represents the mean perception $v_{0}$. $v_{0}$ may differ from the actual value $v_{i}$ to capture potential biases in perception. The variance $\sigma_{0}^{2}$, on the other hand, reflects the level of uncertainty in individuals’ perceptions: all else being equal, a larger variance indicates greater uncertainty about their expected representative utility, making individuals more influenced by incoming information. Therefore, at period 00, the individual’s mean expectation $E[v_{i}|I_{i0}]$ and the variance of such mean expectation $\sigma_{i0}^{2}$ can be expressed as:
@@
Consistent with prior Bayesian learning studies and the concept of rational expectations, the mean of the signal is assumed to equal the true representative utility (Zhang et al. 2020, Huang et al. 2014). Regarding the signal variance, similar to any Bayesian updating process, it influences the relative weight given to prior beliefs versus the incoming signal when constructing the posterior belief, thereby reflecting individuals’ learning speed in converging towards the actual representative utility (Ho et al. 2017, Erdem and Keane 1996). As introduced by our RQ1 (learning divide), individuals with different backgrounds may exhibit heterogeneous learning abilities. To explore such potential disparities, we link observable attributes to the variance as specified in Equation 9. Specifically, within this equation, we apply an exponential transformation to a linear combination of attributes to ensure that the variance $\sigma_{s_{i}}^{2}$ remains positive and covers the entire positive domain.
$\sigma_{s_{i}}^{2}=\exp(\gamma_{0i}+\gamma_{1}\cdot HighEdu_{i}+\gamma_{2}\cdot Age_{i}+\gamma_{3}\cdot Male_{i}+\gamma_{4}\cdot White_{i}+\gamma_{5}\cdot IT_{i})$ (9)
@@
We now further illustrate the perception updating process based on signals. To improve estimation efficiency, we aggregate data and analyze users’ learning processes on a daily basis (Zhou et al. 2022). At the beginning of day $(t-1)$, an individual’s perceived representative utility of ChatGPT is characterized by a normal distribution with mean $E[v_{i}|I_{i(t-1)}]$ and variance $\sigma_{i(t-1)}^{2}$. Throughout day $(t-1)$, user $i$ interacts with ChatGPT for $W_{i(t-1)s}$ times and accesses $W_{i(t-1)n}$ related news articles. Using these signals, the user updates their belief according to the Bayesian rule (DeGroot 2005, Erdem and Keane 1996), forming a new belief at the end of the day that the actual representative utility is drawn from the distribution $N(E[v_{i}|I_{it}],\sigma_{it}^{2})$ where
$E[v_{i}|I_{it}]$ $=\frac{\frac{E[v_{i}|I_{i(t-1)}]}{\sigma_{i(t-1)}^{2}}+\sum_{k=1}^{W_{i(t-1)s}}\frac{s_{i(t-1)j}}{\sigma_{s_{i}}^{2}}+\sum_{j=1}^{W_{i(t-1)n}}\frac{n_{i(t-1)j}}{\sigma_{n}^{2}}}{{\frac{1}{\sigma_{i(t-1)}^{2}}+\frac{W_{i(t-1)s}}{\sigma_{s_{i}}^{2}}}+\frac{W_{i(t-1)n}}{\sigma_{n}^{2}}}$ (10)
@@
As additional signals are received, the perceived representative utility $E[v_{i}|I_{it}]$ gradually converges towards the true representative utility $v_{i1}$, albeit with some noise, while the associated uncertainty $\sigma_{it}^{2}$ progressively diminishes towards zero.
@@ 4.3 Unobserved Heterogeneity and Likelihood Function
In this binary choice scenario, each individual selects the option that maximizes their expected utility by comparing $E[u_{it1}|I_{it}]$ and $E[u_{it0}|I_{it}]$. Assuming that the unobserved utility shock follows a Type-1 extreme value distribution, the difference between these utilities is characterized by a logistic distribution (Train 2009). This assumption enables a closed-form expression for the choice probability of selecting ChatGPT, based on the representative perceived utilities ($E[v_{i}|I_{it}]$ and $c$), using the logit model:
$Pr(d_{itw}=1)$ $=\frac{\exp(E[v_{i}|I_{it}])}{\exp(E[v_{i}|I_{it}])+\exp(c)}$ (12)
@@
$l_{it}(W_{itd},W_{its}\mid\bm{\theta})$ $=\left(\frac{\exp(E[v_{i}|I_{it}])}{\exp(E[v_{i}|I_{it}])+\exp(c)}\right)^{W_{its}}\quad\times\left(\frac{\exp(c)}{\exp(E[v_{i}|I_{it}])+\exp(c)}\right)^{W_{itd}-W_{its}}$ (13)
where $\bm{\theta}$ represents the parameter vector governing the individual’s utility function or learning process: $\bm{\theta}=[c,\alpha_{0i},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\gamma_{0i},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\sigma_{n}^{2}]$.
Throughout the model setup, heterogeneity in both the utility function and the signal function is captured using observable variables and is assumed to be independent, conditional on these observables. We now relax this assumption by employing a latent class approach (Kamakura and Russell 1989). Specifically, each individual is assumed to have a probability of belonging to a particular latent class, which introduces two class-specific constants related to the utility function and the signal function, respectively. This approach effectively captures and accounts for heterogeneity that is not reflected in the observable variables. Additionally, it allows for unobserved dependencies. For instance, a negative correlation between utility gain and learning speed implies that, all else being equal, individuals who learn more quickly (i.e., have lower variance) also tend to derive relatively higher utility gains from using ChatGPT.
In practice, for latent class 1, we assign the probability of individuals belonging to this class as $Pr(Class_{i}=1)=\frac{1}{1+\exp(\lambda)}$, with associated parameters $\alpha_{0i}=\alpha_{0}$ and $\gamma_{0i}=\gamma_{0}$. For latent class 2, the probability is given by $Pr(Class_{i}=2)=\frac{\exp(\lambda)}{1+\exp(\lambda)}$, with associated parameters $\alpha_{0i}=\alpha_{0}+\Delta\alpha_{0}$ and $\gamma_{0i}=\gamma_{0}+\Delta\gamma_{0}$.
@@
where $\bm{\theta^{\prime}}$ includes $[c,\alpha_{0},\Delta\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\gamma_{0},\Delta\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\sigma_{n}^{2},\lambda]$.
@@ 5.1 Model Estimation
$L(\bm{\theta^{\prime}})=\prod_{i=1}^{I}\prod_{t=1}^{T_{i}}l_{it}(W_{itd},W_{its}\mid\bm{\theta^{\prime}})$ (15)
A common approach to estimate the parameters of interest $\bm{\theta^{\prime}}$ is to find those values that maximize the empirical likelihood (Train 2009). However, this approach is not straightforward for the Bayesian learning model given the available data: we can only observe each individual’s attributes such as $HighEdu_{i}$, $Age_{i}$, $Male_{i}$, $White_{i}$, $IT_{i}$, and their usage of ChatGPT ($W_{its}$) and traditional search engines ($W_{itd}-W_{its}$) during each period. We aim to derive the unknown $\bm{\theta^{\prime}}$, but we cannot directly observe users’ perceived representative utility for ChatGPT ($E[v_{i}|I_{it}]$) at each period or the stochastic error terms ($\varepsilon_{it0}$, $\varepsilon_{it1}$). Similar to the mixed logit model, slight modifications can render a fully closed-form expression of the likelihood unfeasible (Train 2009). Although there is a direct mathematical relationship between perceived utility and choice probability under the logit model, deriving the parameters of interest from perceived utility lacks a closed-form solution, as we cannot observe the initial belief distribution $N(v_{0},\sigma_{0}^{2})$ or the exact signals ($s_{itk}$ and $n_{itk}$) (Erdem and Keane 1996).
@@ 5.2 Identification
In the Bayesian learning process, users learn the true representative utility from signals and gradually update their mean belief from $v_{0}$ to $v_{i}$. Following prior literature (Wang and Yang 2024, Ching et al. 2013), $v_{0}$ and $v_{i}$ cannot be simultaneously identified, and therefore $v_{0}$ is normalized to zero to ensure model identifiability. With $v_{0}$ held constant, the outside option’s representative utility $c$ is identified based on users’ initial choice probabilities between ChatGPT and traditional search engines. A decrease in $c$ corresponds to a relative increase in the net utility from using ChatGPT, thereby increasing the likelihood of ChatGPT adoption in the absence of other information.
Leveraging $c$, the representative utility of ChatGPT $v_{i}$ is identified based on users’ steady-state choice probabilities. The utility parameters ($\alpha_{0i},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}$) are then derived based on the heterogeneous steady-state choice probabilities among individuals with different backgrounds. For example, an increase in $\alpha_{1}$ implies a greater impact of education on representative utility, making higher-educated users more likely to choose ChatGPT over traditional search engines.
Furthermore, the variance of the signals ($\sigma_{is}^{2}$ and $\sigma_{n}^{2}$) is identified through the variation in users’ belief updates after receiving different signals. As with the identification of the mean value, we cannot simultaneously identify the prior variance and the signal variance. Thus, the prior variance $\sigma_{0}^{2}$ is fixed at $\exp(4)$ to reflect substantial uncertainty that individuals face when they have no prior information about ChatGPT. The core findings are robust to different choices of prior variance values. As individuals receive varying numbers of news signals over time, differences in how their beliefs evolve provide information about $\sigma_{n}^{2}$. Similarly, the signal parameters ($\gamma_{0i},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5}$) are identified based on the variation in learning speeds across individuals with different backgrounds.
Finally, since individuals exhibit different dynamics depending on their latent class (Zheng et al. 2020), the parameter $\lambda$ that governs the class probability, along with the associated class-specific parameters ($\alpha_{0}$, $\Delta\alpha_{0}$, $\gamma_{0}$, and $\Delta\gamma_{0}$), can be identified. Prior to estimating the model with empirical data, a Monte Carlo simulation is conducted to ensure that all parameters can be accurately identified within the proposed model structure and estimation procedure.

### v3

@@ 5 Model
Motivated by our theoretical inquiry and exploratory analyses, we develop a Bayesian learning model to investigate the dynamics of adoption behaviors. Whenever a user would like to access resources from the Internet, she must decide whether to start the process with ChatGPT or traditional search engines. This decision is based on the perceived utility derived from each option, which evolves through a learning process influenced by the user’s prior experiences. Consequently, the final choice is determined through utility maximization. The utility function, learning process, and decision-making procedure will be elaborated upon in the following subsections.
@@ 5.1 Utility Function
Given that the traditional search engine has been active for a long period, we use it as the baseline and assume that its utility $u_{it0}$ can be decomposed into a constant term $c_{i}$ and a stochastic error term $\varepsilon_{it0}$, in line with the Bayesian learning literature (Zheng et al. 2020):
$u_{it0}=c_{i}+\varepsilon_{it0}$ (1)
@@
As introduced in RQ2, users with different social characteristics may derive varying levels of utility from using ChatGPT. Therefore, we model each individual’s attributes as contributing linearly to her representative utility, allowing us to characterize the potential utility divide.
$v_{i}=$ $\ \alpha_{0i}+\alpha_{1}\cdot College_{i}+\alpha_{2}\cdot AgeHigh_{i}+\alpha_{3}\cdot AgeMid_{i}+\alpha_{4}\cdot Male_{i}$
$\ +\alpha_{5}\cdot White_{i}+\alpha_{6}\cdot LLMOE_{i}+\alpha_{7}\cdot English_{i}$ (3)
Note that in other discrete choice settings (Train 2009), only differences in utility are relevant; therefore, only one constant term is needed in the utility functions of the two alternatives to capture the baseline relative differences between ChatGPT and the traditional search engine. However, in our learning model, where we normalize the distribution of individuals’ initial beliefs, both constants can be identified, as discussed in Section 6. Additionally, the parameters $(\alpha_{1},\ldots,\alpha_{7})$ also capture differences in relative utility instead of absolute utility. For instance, a positive $\alpha_{1}$ would indicate that users with higher education tend to derive greater relative gains from using ChatGPT compared to traditional search engines.
Since ChatGPT has been launched only recently, individuals may not initially be fully certain about ChatGPT’s utility $u_{it1}$. Therefore, before a user $i$ decides whether to use ChatGPT to fulfill her query, she first forms an expectation of the utility based on the information available up to period $t$ (hereinafter referred to as “perceived utility”), and then choose the alternative that maximizes her expected utility. The perception formation process follows the Bayesian learning framework (Wang and Yang 2024), and will be elucidated in Section 5.2.
@@ 5.2 Individuals’ Learning Process
Since traditional search engines have been available for an extended period, individuals are already familiar with them, and their beliefs are assumed to have stabilized prior to the observation period. Consequently, similar to other Bayesian learning models (Zheng et al. 2020), the perceived representative utility of the outside option $E[c_{i}|I_{it}]$ is assumed to be constant for consumers a priori and is not influenced by new information.
$E[u_{it0}|I_{it}]=E[c_{i}|I_{it}]+\varepsilon_{it0}=c_{i}+\varepsilon_{it0}$ (6)
In contrast, the utility of the newly introduced ChatGPT is less certain to users, resulting in a dynamic update of their belief $E[u_{it1}|I_{it}]$ over time as they gather more information and experience with the tool. Following the Bayesian learning literature (Wang and Yang 2024, Erdem and Keane 1996), we assume that prior to any direct experience with ChatGPT, individuals possess some prior information regarding the representative utility of using ChatGPT, which is modeled as following a normal distribution $N(v_{0},\sigma_{0}^{2})$. The individual perceives that the actual representative utility $v_{i}$ is drawn from this distribution, where the mean of this distribution represents the mean perception $v_{0}$. $v_{0}$ may differ from the actual value $v_{i}$, which can reflect the initial skepticism about ChatGPT’s representative utility. Users who are more skeptical may begin with a lower $v_{0}$, leading to a larger gap with the true representative utility $v_{0}$ and requiring a longer belief-updating process to converge toward the true representative utility. The variance $\sigma_{0}^{2}$, on the other hand, reflects the level of uncertainty in individuals’ perceptions: all else being equal, a larger variance indicates greater uncertainty about their expected representative utility, making individuals more influenced by incoming information. Therefore, at period 0, the individual’s mean expectation $E[v_{i}|I_{i0}]$ and the variance of such mean expectation $\sigma_{i0}^{2}$ can be expressed as:
@@
Consistent with prior Bayesian learning studies and the concept of rational expectations, the mean of the signal is assumed to equal the true representative utility (Zhang et al. 2020, Huang et al. 2014). Regarding the signal variance, similar to any Bayesian updating process, it influences the relative weight given to prior beliefs versus the incoming signal when constructing the posterior belief, thereby reflecting individuals’ learning speed in converging towards the actual representative utility (Ho et al. 2017, Erdem and Keane 1996). As introduced by our RQ1 (learning divide), individuals with different backgrounds may exhibit heterogeneous learning abilities. To explore such potential disparities, we link observable attributes to the variance as specified in Equation 5.2. Specifically, within this equation, we apply an exponential transformation to a linear combination of attributes to ensure that the variance $\sigma_{s_{i}}^{2}$ remains positive and covers the entire positive domain.
$\sigma_{s_{i}}^{2}=$ $\ \exp(\gamma_{0i}+\gamma_{1}\cdot College_{i}+\gamma_{2}\cdot AgeHigh_{i}+\gamma_{3}\cdot AgeMid_{i}+\gamma_{4}\cdot Male_{i}$
$\ +\gamma_{5}\cdot White_{i}+\gamma_{6}\cdot LLMOE_{i}+\gamma_{7}\cdot English_{i})$ (9)
@@
We now further illustrate the perception updating process based on signals. To improve estimation efficiency, we aggregate data and analyze users’ learning processes on a daily basis (Zhou et al. 2022). At the beginning of day (t1)(t-1), an individual’s perceived representative utility of ChatGPT is characterized by a normal distribution with mean $E[v_{i}|I_{i(t-1)}]$ and variance $\sigma_{i(t-1)}^{2}$. Throughout day (t1)(t-1), user $i$ interacts with ChatGPT for $W_{i(t-1)s}$ times and accesses $W_{i(t-1)n}$ related news articles. Using these signals, the user updates their belief according to the Bayesian rule (DeGroot 2005, Erdem and Keane 1996), forming a new belief at the end of the day that the actual representative utility is drawn from the distribution $N(E[v_{i}|I_{it}],\sigma_{it}^{2})$ where
$E[v_{i}|I_{it}]$ $=\frac{\frac{E[v_{i}|I_{i(t-1)}]}{\sigma_{i(t-1)}^{2}}+\sum_{k=1}^{W_{i(t-1)s}}\frac{s_{i(t-1)k}}{\sigma_{s_{i}}^{2}}+\sum_{j=1}^{W_{i(t-1)n}}\frac{n_{i(t-1)j}}{\sigma_{n}^{2}}}{{\frac{1}{\sigma_{i(t-1)}^{2}}+\frac{W_{i(t-1)s}}{\sigma_{s_{i}}^{2}}}+\frac{W_{i(t-1)n}}{\sigma_{n}^{2}}}$ (10)
@@
As additional signals are received, the perceived representative utility $E[v_{i}|I_{it}]$ gradually converges towards the true representative utility $v_{i}$, albeit with some noise, while the associated uncertainty $\sigma_{it}^{2}$ progressively diminishes towards zero.
@@ 5.3 Unobserved Heterogeneity and Likelihood Function
In this binary choice scenario, each individual selects the option that maximizes their expected utility by comparing $E[u_{it1}|I_{it}]$ and $E[u_{it0}|I_{it}]$. Assuming that the unobserved utility shock follows a Type-1 extreme value distribution, the difference between these utilities is characterized by a logistic distribution (Train 2009). This assumption enables a closed-form expression for the choice probability of selecting ChatGPT, based on the representative perceived utilities ($E[v_{i}|I_{it}]$ and $c_{i}$), using the logit model:
$Pr(d_{itw}=1)$ $=\frac{\exp(E[v_{i}|I_{it}])}{\exp(E[v_{i}|I_{it}])+\exp(c_{i})}$ (12)
@@
$l_{it}(W_{itd},W_{its}\mid\bm{\theta})$ $=\left(\frac{\exp(E[v_{i}|I_{it}])}{\exp(E[v_{i}|I_{it}])+\exp(c_{i})}\right)^{W_{its}}\quad\times\left(\frac{\exp(c_{i})}{\exp(E[v_{i}|I_{it}])+\exp(c_{i})}\right)^{W_{itd}-W_{its}}$ (13)
where $\bm{\theta}$ represents the parameter vector governing the individual’s utility function or learning process: $\bm{\theta}=[c_{i},\alpha_{0i},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\gamma_{0i},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\sigma_{n}^{2}]$.
Throughout the model setup, heterogeneity in both the utility function and the signal function is captured using observable variables and is assumed to be independent, conditional on these observables. We now relax this assumption by employing a latent class approach (Kamakura and Russell 1989). Specifically, each individual is assumed to have a probability of belonging to a particular latent class, which introduces class-specific constants related to the utility function and the signal function, respectively. This approach effectively captures and accounts for heterogeneity that is not reflected in the observable variables. Additionally, it allows for unobserved dependencies. For instance, a negative correlation between utility gain and learning speed implies that, all else being equal, individuals who learn more quickly (i.e., have lower variance) also tend to derive relatively higher utility gains from using ChatGPT.
We compare models with varying numbers of latent classes using standard model selection criteria and ultimately select a two-class specification. Without loss of generality, we present the model formulation for the case of two latent classes and report the model selection results in Appendix R. Specifically, for latent class 1, we assign the probability of individuals belonging to this class as $Pr(Class_{i}=1)=\frac{1}{1+\exp(\lambda)}$, with associated parameters $c_{i}=c$, $\alpha_{0i}=\alpha_{0}$, and $\gamma_{0i}=\gamma_{0}$. For latent class 2, the probability is given by $Pr(Class_{i}=2)=\frac{\exp(\lambda)}{1+\exp(\lambda)}$, with associated parameters $c_{i}=c+\Delta c$, $\alpha_{0i}=\alpha_{0}+\Delta\alpha_{0}$ and $\gamma_{0i}=\gamma_{0}+\Delta\gamma_{0}$.
@@
where $\bm{\theta^{\prime}}$ includes $c,\Delta c,\alpha_{0},\Delta\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\gamma_{0},\Delta\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\sigma_{n}^{2},\lambda]$.
@@ 6.1 Model Estimation
$L(\bm{\theta^{\prime}})=\prod_{i=0}^{I-1}\prod_{t=0}^{T_{i}-1}l_{it}(W_{itd},W_{its}\mid\bm{\theta^{\prime}})$ (15)
A common approach to estimate the parameters of interest $\bm{\theta^{\prime}}$ is to find those values that maximize the empirical likelihood (Train 2009). However, this approach is not straightforward for the Bayesian learning model given the available data: we can only observe each individual’s attributes such as $College_{i}$, $AgeHigh_{i}$, $AgeMid_{i}$, $Male_{i}$, $White_{i}$, $LLMOE_{i}$, $English_{i}$, and their usage of ChatGPT ($W_{its}$) and traditional search engines ($W_{itd}-W_{its}$) during each period. We aim to derive the unknown $\bm{\theta^{\prime}}$, but we cannot directly observe users’ perceived representative utility for ChatGPT ($E[v_{i}|I_{it}]$) at each period or the stochastic error terms ($\varepsilon_{it0}$, $\varepsilon_{it1}$). Similar to the mixed logit model, slight modifications can render a fully closed-form expression of the likelihood unfeasible (Train 2009). Although there is a direct mathematical relationship between perceived utility and choice probability under the logit model, deriving the parameters of interest from perceived utility lacks a closed-form solution, as we cannot observe the initial belief distribution $N(v_{0},\sigma_{0}^{2})$ or the exact signals ($s_{itk}$ and $n_{itk}$) (Erdem and Keane 1996).
@@ 6.2 Identification
In the Bayesian learning process, users learn the true representative utility from signals and gradually update their mean belief from $v_{0}$ to $v_{i}$. Following prior literature (Wang and Yang 2024, Ching et al. 2013), $v_{0}$ and $v_{i}$ cannot be simultaneously identified, and therefore $v_{0}$ is normalized to zero to ensure model identifiability. With $v_{0}$ held constant, the outside option’s representative utility $c_{i}$ is identified based on users’ initial choice probabilities between ChatGPT and traditional search engines. A decrease in $c_{i}$ corresponds to a relative increase in the net utility from using ChatGPT, thereby increasing the likelihood of ChatGPT adoption in the absence of other information.
Leveraging $c_{i}$, the representative utility of ChatGPT $v_{i}$ is identified based on users’ steady-state choice probabilities. The utility parameters ($\alpha_{0i},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{6},\alpha_{7}$) are then derived based on the heterogeneous steady-state choice probabilities among individuals with different backgrounds. For example, an increase in $\alpha_{1}$ implies a greater impact of education on representative utility, making users who receive college education more likely to choose ChatGPT over traditional search engines.
Furthermore, the variance of the signals ($\sigma_{s_{i}}^{2}$ and $\sigma_{n}^{2}$) is identified through the variation in users’ belief updates after receiving different signals. As with the identification of the mean value, we cannot simultaneously identify the prior variance and the signal variance. Thus, the prior variance $\sigma_{0}^{2}$ is fixed at $\exp(4)$ to reflect substantial uncertainty that individuals face when they have no prior information about ChatGPT. The core findings are robust to different choices of prior variance values as shown in Appendix U. As individuals receive varying numbers of news signals over time, differences in how their beliefs evolve provide information about $\sigma_{n}^{2}$. Similarly, the signal parameters ($\gamma_{0i},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\gamma_{6},\gamma_{7}$) are identified based on the variation in learning speeds across individuals with different backgrounds.
Finally, since individuals exhibit different dynamics depending on their latent class (Zheng et al. 2020), the parameter $\lambda$ that governs the class probability, along with the associated class-specific parameters ($c$, $\Delta c$, $\alpha_{0}$, $\Delta\alpha_{0}$, $\gamma_{0}$, and $\Delta\gamma_{0}$), can be identified. Prior to estimating the model with empirical data, a Monte Carlo simulation is conducted to ensure that all parameters can be accurately identified within the proposed model structure and estimation procedure.